OpenStudy (anonymous):

What is x in (2x-25)+3x+(x+6)+(4x-11)=360

OpenStudy (precal):

2x-25+3x+x+6+4x-11=360 10x-30=360 add 30 to both sides 10x=390 divide by 10 to both sides x=39

OpenStudy (anonymous):

add like terms... and solve =0

OpenStudy (anonymous):

10x-30=360 10x=390 x=39 XD hey iHelp XD

OpenStudy (anonymous):

hi yama =)

OpenStudy (anonymous):

Thanks